Jessie Co. issued an $8 million face amount of 9%, 30-year bonds on April 1, 2019. The bonds pay interest on an annual basis on March 31 each year. Calculate the interest expense that Jessie Co. will show with respect to these bonds in its income statement for the fiscal year ended September 30, 2019, assuming that the premium of $68,000 is amortized on a straight-line basis.

What is the Accrued interest payable?

Premium amortization?

Interest expense for 6 months?

Can you please show me the steps of how the problem is solved? Thank you so much

Premium amortization
Answer #1

Accrued interest expense = 8000000*9%*6/12 = 360000

Premium amortization = 68000/60 = 1133

Interest expense = 360000-1133 = 358867